A rectangle is drawn so that the width is 2 feet shorter than the length. The area of the rectangle is 3 square feet. Find the length of the rectangle.

Answer by Theo(13342)

area = length * width = 3
length = x and width = x - 2
area = x * (x - 2) = x^2 - 2x = 3
subtract 3 from both sides of that equation to get x^2 - 2x - 3 = 0.
factor to get (x - 3) * (x + 1) = 0
solve for x to get x = 3 or x = -1
x has to be positive, so x = 3
length = 3
width = 1
length = 3 is your solution.